OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (5, 15) and (9, 2) B. (15, 5) and (2, 9) C. (5, 10) and (2, 9) D. (5, 15) and (2, 9) GIVING MEDALS

OpenStudy (blake57roger):

d

OpenStudy (anonymous):

could you explain it please

OpenStudy (anonymous):

no

OpenStudy (blake57roger):

1.) y = 2x+5 y = 2(8)+5 y = 21 Solution: (8,21) 2.) y = 2x+5 y = 2(-1)+5 y = 3 Solution: (-1,3) 3.) y = 2x+5 -1 = 2x+5 -6 = 2x x = -3 Solution: (-3,-1) Solutions: y = 2x+5; (8,21), (-1,3), (-3,-1)

OpenStudy (blake57roger):

does that help

OpenStudy (anonymous):

yes thank you :) heres your medal

OpenStudy (blake57roger):

thanks any more

OpenStudy (anonymous):

possibly ill message you if i need help thanks :)

OpenStudy (blake57roger):

k